# Bridging Fairness and Explainability: Can Input-Based Explanations Promote Fairness in Hate Speech Detection?

### v1

@@ Appendix B Model Performance and Fairness Results
Llama3.2 Zero-shot 63.78/74.62/71.27 1.45/2.35/24.67 11.03/3.52/36.81 10.54/1.03/2.95 2.13/2.94/3.83
Few-shot 46.45/28.55/42.23 1.20/0.90/21.27 3.87/1.82/30.69 0.80/0.19/1.66 0.11/0.13/0.33
@@
Qwen3 Zero-shot 66.75/77.25/77.33 3.50/3.75/16.33 4.21/3.78/17.40 0.80/4.05/5.89 3.05/2.31/3.67
Few-shot 56.12/61.44/75.67 7.75/1.63/9.67 9.95/2.11/11.48 0.23/1.69/2.44 3.93/4.96/4.13
@@
Llama3.2 Zero-shot 54.00/70.50/65.17 8.50/1.00/25.67 10.91/1.53/31.87 0.20/4.56/8.33 2.39/3.00/4.28
Few-shot 27.62/23.69/83.00 4.75/4.37/6.00 6.92/4.90/5.74 2.42/3.74/28.13 0.04/0.03/0.02
Tables 4 and 5 show the task performance and fairness scores for the default / zero-shot and debiased models on the Civil Comments and Jigsaw datasets respectively. To better identify the differences between different debiasing methods, we conduct an analysis based on how often a debiasing method successfully reduces the average individual unfairness ($\text{Avg}_{\text{iu}}$) and maintains the task performance (Accuracy) of the default / zero-shot model.
@@
We find that the debiasing methods (fairness imagination and fairness instruction) for the decoder-only models consistently improve the task performance across all bias types and datasets. Contrary to this, we see increases in average individual unfairness of the fairness imagination approach for race and gender bias across both datasets. Only for religion, fairness imagination leads to a decrease of the individual unfairness. For fairness instruction, we observe a consistent improvement across all three bias types and both datasets, showing the clear superiority of the approach. The consistency of the results is especially surprising when considering that both decoder-only models are instruction-tuned, and that Chen et al. [2025] identify a bias amplification effect from instruction tuning. We conclude that fairness instruction is a good baseline to evaluate other debiasing methods for decoder-only models.
Appendix C Bias Detection Results
@@
We present the full fairness correlation results of encoder- and decoder-only models with different debiasing methods on Civil Comments and Jigsaw in Figures6, 7, 8, 9. Consistent with findings presented in the main text, Occlusion- and L2-based explanation methods achieve strong fairness correlations across different setups.
@@
Appendix D Model Selection Results
Explanation-Based Metrics
@@ Appendix D Model Selection Results
  • Average absolute sensitive token reliancee: used to predict average individual unfairness, under the assumption that higher reliance on sensitive tokens implies greater sensitivity to group substitutions.
@@
Among these, only average absolute sensitive token reliance exhibits rank correlations above random chance with its target fairness metric (individual unfairness). The correlations for other metrics remain at chance level. Figures 10, 11, 12, 13 demonstrate that no explanation methods can consistently match baseline rank correlation results. Figures 14, 15, 16, 17 further reveal that explanation methods are not able to select the fairest models. These findings underline the unreliability of explanation-based model selection.
@@
Figure 10: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher correlation values indicate greater effectiveness in ranking models. All indicates the model is trained on all bias types.
@@
Figure 11: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher correlation values indicate greater effectiveness in ranking models. All indicates the model is trained on all bias types.
@@
Figure 12: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set on Jigsaw. The validation set size is 200. Higher correlation values indicate greater effectiveness in ranking models. All indicates the model is trained on all bias types.
@@
Figure 13: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set on Jigsaw. The validation set size is 200. Higher correlation values indicate greater effectiveness in ranking models. All indicates the model is trained on all bias types.
@@
Figure 14: MRR@1 results on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models. All indicates the model is trained on all bias types.
@@
Figure 15: MRR@1 results on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models. All indicates the model is trained on all bias types.
@@
Figure 16: MRR@1 results on Jigsaw. The validation set size is 200. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models. All indicates the model is trained on all bias types.
@@
Figure 17: MRR@1 results on Jigsaw. The validation set size is 200. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models. All indicates the model is trained on all bias types.
Appendix E Bias Mitigation Results
@@
The complete bias mitigation results are presented in Figures 18, 19, 20, 21. The findings are in line with conclusions from the main paper, that explanation-based debiasing can effectively reduce model biases across different fairness metrics, bias types, models, and datasets. In addition, the accuracy-fairness harmonic mean results shown in Figures 222324, 25 demonstrate that explanation-based debiasing achieves comparable or superior balance between fairness and task performance than default models and traditional debiasing approaches.
@@
Appendix F Fairness Correlations in Explanation-Debiased Models

### v2

@@ Appendix D Model Performance and Fairness Results
Llama3.2-3B-Instruct Zero-shot 63.78/74.62/71.27 1.45/2.35/24.67 11.03/3.52/36.81 10.54/1.03/2.95 2.13/2.94/3.83
Few-shot 67.80/79.80/80.10 1.60/1.70/18.20 2.49/0.08/6.73 6.73/6.05/10.77 1.39/2.05/1.90
@@
Qwen3-4B Zero-shot 69.55/79.75/77.50 0.60/0.00/17.40 7.13/1.40/21.07 13.25/3.71/5.17 2.55/2.41/3.32
Few-shot 70.15/80.73/79.53 1.80/0.65/18.93 10.02/2.50/19.31 11.89/9.15/5.57 3.18/3.34/3.76
Qwen3-8B Zero-shot 59.27/69.23/66.30 1.25/0.15/26.80 8.18/0.07/42.05 4.65/0.80/3.02 3.27/3.40/4.74
Few-shot 66.97/77.30/77.47 0.05/0.00/23.27 6.14/2.73/29.51 7.95/7.64/2.34 4.23/4.58/5.96
Fairness imagination 62.10/72.92/69.97 1.60/0.55/21.87 7.80/2.62/32.27 4.28/5.42/9.43 2.54/2.08/2.58
Fairness instruction 66.50/75.15/73.90 0.90/0.10/21.20 8.03/1.76/28.60 8.79/4.59/7.45 2.45/2.95/3.15
@@
Llama3.2-3B-Instruct Zero-shot 54.00/70.50/65.17 8.50/1.00/25.67 10.91/1.53/31.87 0.20/4.56/8.33 2.39/3.00/4.28
Few-shot 73.12/88.62/86.83 7.25/0.50/7.67 13.01/0.16/4.83 15.05/9.08/23.17 1.63/1.68/2.00
Qwen3-4B Zero-shot 66.75/77.25/77.33 3.50/3.75/16.33 4.21/3.78/17.40 0.80/4.05/5.89 3.05/2.31/3.67
Few-shot 57.88/68.06/77.83 8.75/2.12/9.33 11.52/1.80/10.86 1.49/5.79/9.45 3.60/4.31/4.18
@@
Qwen3-8B Zero-shot 48.12/59.50/56.50 7.25/0.00/13.00 9.68/0.37/18.59 1.31/4.92/6.30 3.31/3.47/5.52
Few-shot 53.75/67.19/77.17 5.50/1.12/8.67 6.18/1.53/10.51 3.41/1.95/4.88 4.50/5.04/5.99
Fairness imagination 51.62/67.50/61.83 4.25/0.75/9.67 5.24/0.56/11.23 1.05/3.49/6.45 2.51/2.02/2.55
Fairness instruction 60.25/71.19/67.50 8.50/1.87/12.00 10.57/2.23/14.22 0.93/1.90/2.10 2.46/3.13/3.60
Tables 5 and 6 show the task performance and fairness scores for the default/zero-shot and debiased models on the Civil Comments and Jigsaw datasets respectively. To better identify the differences between different debiasing methods, we conduct an analysis based on how often a debiasing method successfully reduces the average individual unfairness ($\text{Avg}_{\text{iu}}$) and maintains the task performance (Accuracy) of the default/zero-shot model.
@@
We find that the debiasing methods (fairness imagination and fairness instruction) for the decoder-only models consistently improve the task performance across all bias types and datasets. Contrary to this, we see increases in average individual unfairness of the fairness imagination approach for race and gender bias on Llama3.2-3B-Instruct and Qwen3-4B across both datasets. Only for religion, fairness imagination leads to a consistent decrease of the individual unfairness across models. For fairness instruction, we observe a consistent improvement across all three bias types and both datasets, showing the clear superiority of the approach. The consistency of the results is especially surprising when considering that both decoder-only models are instruction-tuned and aligned with human values, and that Chen et al. (2025) identify a bias amplification effect from instruction tuning. We conclude that fairness instruction is a good baseline to evaluate other debiasing methods for decoder-only models.
Appendix E Bias Detection Results
@@
We present the full fairness correlation results of encoder- and decoder-only models with different debiasing methods on Civil Comments and Jigsaw in Figures7, 8, 9, 10. Consistent with findings presented in the main text, Occlusion- and L2-based explanation methods achieve strong fairness correlations across different setups.
@@
Appendix F Model Selection Results
Explanation-Based Metrics and Fair Model Selection Results
@@ Appendix F Model Selection Results
  • Average absolute sensitive token reliance: used to predict average individual unfairness, under the assumption that higher reliance on sensitive tokens implies greater sensitivity to group substitutions.
@@
Among these, only average absolute sensitive token reliance exhibits rank correlations above random chance with its target fairness metric (individual unfairness). The correlations for other metrics remain at chance level. Figures 11, 12, 13, 14 demonstrate that no explanation methods can consistently match baseline rank correlation results.
Figures 15, 16, 17, 18 further reveal that explanation methods are not able to robustly select the fairest models. These findings underline the unreliability of explanation-based model selection.
@@
Figure 11: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set for encoder-only models on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher correlation values indicate greater effectiveness in ranking models. All indicates the model is trained on all bias types.
@@
Figure 12: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set for decoder-only models on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher correlation values indicate greater effectiveness in ranking models.
@@
Figure 13: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set for encoder-only models on Jigsaw. The validation set size is 200. Higher correlation values indicate greater effectiveness in ranking models. All indicates the model is trained on all bias types.
@@
Figure 14: Rank correlations between validation set average absolute sensitive token reliance and individual unfairness on the test set for decoder-only models on Jigsaw. The validation set size is 200. Higher correlation values indicate greater effectiveness in ranking models.
@@
Figure 15: MRR@1 results for encoder-only models on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models. All indicates the model is trained on all bias types.
@@
Figure 16: MRR@1 results for decoder-only models on Civil Comments. The validation set sizes are 500 for race, 500 for gender, and 200 for religion. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models.
@@
Figure 17: MRR@1 results for encoder-only models on Jigsaw. The validation set size is 200. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models. All indicates the model is trained on all bias types.
@@
Figure 18: MRR@1 results for decoder-only models on Jigsaw. The validation set size is 200. Higher MRR@1 scores indicate explanations are more effective in selecting the fairest models.
Appendix G Bias Mitigation Results
@@
The complete bias mitigation results are presented in Figures 19, 20, 21, 22. The findings are in line with conclusions from the main paper, that explanation-based debiasing can effectively reduce model biases across different fairness metrics, bias types, models, and datasets. In addition, the accuracy-fairness harmonic mean results shown in Figures 232425, 26 demonstrate that explanation-based debiasing achieves comparable or superior balance between fairness and task performance than default models and traditional debiasing approaches.
We additionally report the results of Integrated Gradients for bias mitigation in Table 7. Similar to other explanation methods, IntGrad-based debiasing achieves substantial bias reduction and maintains a good balance between fairness and task performance in $\text{Disp}_{\text{fnr}}$ and $\text{Avg}_{\text{iu}}$.
@@
Table 7: Results of mitigating race bias in BERT models using Intgrad explanations on Civil Comments. For consistency with accuracy, fairness results are reported as $100-\{\text{Disp}_{\text{acc}},\text{Disp}_{\text{fpr}},\text{Disp}_{\text{fnr}},\text{Avg}_{\text{iu}}\}$, so that higher values indicate better debiasing performance. Each column corresponds to models selected by maximizing the fairness-balanced metric with respect to the indicated bias metric. H-Mean indicates harmonic mean between fairness and accuracy. Green (red) indicates the results are better (worse) than the default models.
Method $\text{Disp}_{\text{{acc}}}$ $\text{Disp}_{\text{{fpr}}}$ $\text{Disp}_{\text{{fnr}}}$ $\text{Avg}_{\text{{iu}}}$
Acc Fairness H-Mean Acc Fairness H-Mean Acc Fairness H-Mean Acc Fairness H-Mean
IntGrad L1 78.55 96.66 86.67 78.55 96.66 86.67 77.98 97.86 86.8 78.37 97.02 86.7
IntGrad L2 77.85 96.58 86.21 77.71 96.33 86.02 77.85 96.58 86.21 78.09 97.1 86.56
Default 78.97 98.37 87.61 78.97 98.96 87.84 78.97 91.15 84.62 78.97 96.36 86.8
Appendix H Generalization of Findings
To demonstrate the generalizability of our findings, we present results under additional setups that vary in task, model alignment type, and sensitive token vocabulary. We observe similar results across these setups, suggesting that our findings generalize well beyond the main study conditions.
Task
We evaluated explanation-based bias detection (RQ1) on an additional task, namely sentiment analysis, using the Twitter Sentiment dataset [^9] https://huggingface.co/datasets/shukdevdatta123/twitter_sentiment_preprocessed. Specifically, we selected 1000 gender-related examples (500 referencing males and 500 referencing females) and ran explanation-based bias detection on them. In Figure 27 we report the results on Llama3.2-3B-Instruct and Qwen3-4B models.
Figure 27: Fairness correlation results for race bias on Twitter Sentiment with Llama3.2-3B-Instruct and Qwen3-4B. Higher values indicate that the method is more effective and reliable in detecting biased predictions at inference time.
In Figure 27, we observe patterns in the sentiment analysis task that are similar to those in our main study: certain explanation methods (e.g., occlusion-based and L2-based approaches) can still achieve high fairness correlation scores. This suggests that our findings could extend beyond the hate speech detection task.
Model Alignment Type
We extended our experiments to additional LLMs with different alignment methods. Specifically, we evaluated explanation-based bias detection (RQ1) on two differently aligned LLMs: Llama3.2-3B (pre-trained only, non-instruct, used with few-shot prompting) and Qwen2.5-3B-Instruct (instruction-tuned only). Neither model is aligned to human values, which differs from the models used in our main study. The results are computed for race bias on Civil Comments and shown in Figure 28.
Figure 28: Fairness correlation results for race bias on Civil Comments with Llama3.2-3B and Qwen2.5-3B-Instruct. Both models are differently aligned from models in our main experiments. Higher values indicate that the method is more effective and reliable in detecting biased predictions at inference time.
We observe that certain explanation methods, such as Occlusion, consistently achieve high fairness correlations. This suggests that our findings generalize across LLMs with different alignment settings.
Sensitive Token Vocabulary
In practice, it is often unrealistic to exhaustively enumerate all vocabulary items that may encode sensitive attributes. To assess the applicability of our findings under such conditions, we analyze how varying the coverage of sensitive tokens affects bias detection and mitigation. Specifically, we focus on gender bias and use a small subset of gendered pronouns (“he/his/him” and “she/her”) as sensitive tokens, while computing fairness metrics with the full gender-related vocabulary (222 words per gender). This setup simulates real-world scenarios where the sensitive vocabulary cannot be fully enumerated.
Figure 29: Fairness correlation results when using a reduced sensitive token vocabulary for reliance computation. Results are reported for gender bias on the Civil Comments dataset. Fairness metrics are still computed using the full vocabulary. The reduced vocabulary size has only a marginal effect on fairness correlations.
Figure 30: Fairness and accuracy results for gender bias mitigation with a reduced sensitive token vocabulary. Each column corresponds to models selected by maximizing the fairness–balance metric with respect to the indicated bias metric. Using an incomplete vocabulary yields slightly worse debiasing performance than using the complete vocabulary, while preserving task performance more effectively. Overall, the impact of reduced vocabulary coverage is minimal.
As shown in Figures 29 and 30, reduced vocabulary coverage has minimal impact on explanation-based bias detection and mitigation performance. This result is reassuring, suggesting that explanation methods remain effective in more complex, realistic settings where exhaustive sensitive token coverage is infeasible.
Appendix I Faithfulness as an Indicator of Bias Detection Ability
What factors influence the reliability of explanations in detecting bias? In this section, we examine the relationship between explanation faithfulness and their ability to identify bias, reflected by fairness correlation scores in RQ1. We assess the faithfulness of explanation methods using two perturbation-based metrics: comprehensiveness and sufficiency AOPC (Area Over the Perturbation Curve; DeYoung et al., 2020), computed by masking 5%, 10%, 20%, and 50% of the input tokens. For substitution, we use the [MASK] token in BERT and the [PAD] token in Qwen3-4B. Higher comprehensiveness and lower sufficiency scores indicate more faithful explanations.
Table 8: Faithfulness results of different explanation methods on BERT and Qwen3-4B models.
BERT Qwen3-4B
Attention 4.50 3.20 0.62 10.34 17.20 0.14
Attn rollout 4.37 3.11 0.59 9.04 15.70 0.02
Attn flow 4.01 3.46 0.47 10.57 16.82 0.04
Grad L2 4.82 2.99 0.50 12.30 16.09 0.32
Grad mean 0.77 6.16 0.06 11.41 17.50 0.06
DeepLift L2 4.72 3.09 0.50 12.44 16.17 0.31
DeepLift mean 1.68 5.75 0.06 10.78 18.69 0.06
IxG L2 4.89 2.95 0.49 12.35 16.27 0.32
IxG mean 7.44 1.70 0.30 9.99 18.82 0.17
IntGrad L2 4.81 3.02 0.57 12.33 16.86 0.25
IntGrad mean 10.68 -0.36 0.45 14.21 16.12 0.04
Occlusion 13.16 -0.90 0.62 20.05 13.73 0.27
Occlusion abs 0.79 0.56 0.66 22.48 20.36 0.26
KernelSHAP 5.99 2.30 0.21 11.49 17.86 0.02
DecompX 16.08 -2.77 0.40 - - -
ProgInfer - - - 10.32 17.96 0.025
Appendix J Explanations for Human Fairness Auditing
In addition to evaluating input-based explanations as automatic bias detectors, we also examine their ability to support human auditing of biased model predictions. To this end, we conducted a small-scale human study following the experimental protocol described below.
We randomly sample 48 correctly predicted examples related to race bias from Civil Comments (4*6=24 from BERT and 24 from Qwen3-4B, balanced across all group–class categories). We evaluate six explanation methods: three directed methods (Occlusion, IxG mean, KernelSHAP) and three undirected methods (Occlusion abs, IxG L2, Attention), chosen to cover diverse explanation families and performance characteristics observed in RQ1.
For each example, annotators first read the input text and provide their own toxicity prediction. They are then shown either the three directed explanations or the three undirected explanations for that example. For each explanation, annotators give two ratings on a 1–5 scale: one assessing its interpretability, and another evaluating how much the model’s prediction appears to rely on race-related bias or stereotypes, based on the information conveyed by the explanation.
After annotation, we collect annotators’ perceived bias ratings and measure their correlation with the ground-truth individual unfairness scores. Higher fairness correlation indicates greater effectiveness of an explanation method for human fairness auditing.
Table 9: Fairness correlation of explanation methods for human fairness auditing and their interpretability. Best scores in each explanation type are marked in bold. Higher fairness correlation scores indicate that explanations can better assist humans to detect bias.
Undirected Directed
Attention IxG L2 Occlusion abs KernelSHAP IxG mean Occlusion
Fairness correlation 0.402 0.123 0.433 0.254 -0.078 0.342
Interpretability 2.256 3.179 2.920 2.518 2.439 2.780
Table 10: Fairness correlation of explanation methods for human fairness auditing under different conditions. Higher fairness correlation scores indicate that explanations can better assist humans to detect bias. Green (red) indicates the results are better (worse) than the baseline (all examples).
Undirected Directed
Attention IxG L2 Occlusion abs KernelSHAP IxG mean Occlusion
All examples 0.402 0.123 0.433 0.254 -0.078 0.342
Correct predictions Only 0.572 0.202 0.602 0.217 0.029 0.451
Toxic examples only 0.637 0.118 0.374 0.231 0.134 0.315
High interpretability (score $\geq$ 3) 0.138 0.227 0.288 0.154 -0.043 0.404
Table 9 shows that certain explanation methods, such as Attention, Occlusion, and Occlusion abs, achieve high fairness correlations, suggesting that they can effectively assist humans in detecting bias. Across explanation types, undirected explanations appear more helpful. For example, Occlusion and Occlusion abs produce the same attribution patterns that differ only in directional encoding, yet participants were better able to identify bias using the undirected variant (Occlusion abs). Furthermore, while some methods support both human and automatic bias detection consistently (e.g., Attention and Occlusion abs), others, such as IxG L2, show substantial gaps in performance. This highlights a potential discrepancy between how humans interpret explanations and how our automatic pipeline evaluates them.
We also observe that high interpretability alone does not guarantee better support for human fairness auditing: methods with strong interpretability scores (e.g., IxG L2) still fail to effectively help humans detect bias. Finally, 4 out of 6 annotators reported that undirected explanations helped them detect bias more effectively, noting that they introduce less noise and make annotation easier.
Table 10 further analyzes explanation-assisted human fairness auditing under different conditions. For correctly predicted examples, explanations generally provide stronger support for bias detection. However, the effects of label type and explanation interpretability appear more nuanced and vary across methods. Overall, these results suggest that explanation-assisted human fairness auditing is a promising and interesting direction for future work and warrants further investigation.
Appendix K Hybrid Bias Mitigation Techniques
we conducted preliminary experiments that combine several pre-processing techniques (group balance, group–class balance, and CDA) with an effective explanation-based debiasing method (IxG L1/L2). The resulting individual fairness outcomes, along with comparisons to traditional and explanation-only methods, are presented in Table 11.
Table 11: Each cell shows the $\text{Avg}_{\text{iu}}$ score after applying a combination of pre-processing and explanation-based debiasing methods. Lower values indicate reduced bias. Values in parentheses denote the change relative to using only the corresponding traditional/explanation-based method, where negative values indicate improved debiasing. We observe that hybrid approaches consistently achieve stronger bias mitigation than either method used in isolation.
Race
Group balance Group-class balance CDA
IxG L1 0.012 (-4.492/-1.461) 0.000 (-3.048/-1.473) 0.001 (-0.547/-1.473)
IxG L2 2.162 (-2.342/-0.598) 2.110 (-0.938/-0.650) 0.210 (-0.338/-2.550)
Gender
Group balance Group-class balance CDA
IxG L1 0.005 (-0.594/-0.548) 0.001 (-0.836/-0.552) 0.001 (-0.488/-0.551)
IxG L2 0.308 (-0.291/-0.331) 0.546 (-0.292/-0.093) 0.368 (-0.122/-0.271)
We observe that the hybrid method achieves better bias mitigation effects than using each debiasing method alone, with consistent improvements for both race and gender bias. Based on this, we believe exploring hybrid methods for more effective bias mitigation could be a promising future direction.
Appendix L Fairness Correlations in Explanation-Debiased Models